OpenStudy (anonymous):

HELP PLEASE I AM NEW TO THIS!!!!??? Tom and Jack spend a certain amount of money from their accounts each week at a pet shelter. The table shows the relationship between the amount of money (y) remaining in Tom's account and the number of weeks (x): Function 1: Number of Weeks (x) Amount Remaining (dollars) (y) 1 60 2 53 3 46 4 39 The equation shows the relationship between the amount of money, y, remaining in Jack's account and the number of weeks, x: Function 2: y = –8x + 60 Which statements states and explains which function shows a greater rate of change? F

OpenStudy (anonymous):

@priyar

OpenStudy (anonymous):

@rebeccaxhawaii

OpenStudy (priyar):

rate of change = slope =( y2-y2)/(x2-x1) u know this ? @kalua891

OpenStudy (anonymous):

kind of @priyar

OpenStudy (anonymous):

i forgot to put this Which statements states and explains which function shows a greater rate of change? Function 1, because Tom spends $7 each week and Jack spends –$8 each week Function 1, because Tom spends $21 each week and Jack spends $52 each week Function 2, because Jack spends $60 each week and Tom spends $21 each week Function 2, because Jack spends $8 each week and Tom spends $7 each week

OpenStudy (priyar):

ok take two values y1 and y2 and the corresspondind x1, x2 values and plug them in the abv formula of slope.ok?

OpenStudy (anonymous):

what do you mean? @priyar

OpenStudy (anonymous):

i think it is b function 2

OpenStudy (priyar):

how can u say that? did u solve the Q or just guessing?

OpenStudy (anonymous):

i solved it i put in the numbers did i do something wrong maybe?

OpenStudy (anonymous):

my friend told me it was b but i don't think so now

OpenStudy (priyar):

see pls understand how to solve the problem..otherwise how will u solve other problems like these?

OpenStudy (anonymous):

true

OpenStudy (anonymous):

how can i plug the exponents tho I'm confused? @priyar

OpenStudy (priyar):

its actually simple..! first, y1 and y2 are two value from the right side of the table (since they are denoted by the letter 'y'. thats what the Q says right? :) ) similar explanation for x1 and x2.. any doubts so far?

OpenStudy (anonymous):

nop

OpenStudy (anonymous):

so after that? @priyar

OpenStudy (priyar):

good! so pick out any two values of y and corresponding values of x.. now name them as y1,y2 and x1,x2...what are ur values ? (write them along with their names..)

OpenStudy (anonymous):

y 60 and y 46?

OpenStudy (anonymous):

@priyar

OpenStudy (anonymous):

i really have to go!? @priyar

OpenStudy (priyar):

sry.. had connectivity issues..@kalua891 no i meant values next to each other..only then we c an determine the rate of change..right? so i have y1=53 x1=2 y2=60 x2=1 Now plug them in..

OpenStudy (priyar):

@kalua891?

OpenStudy (priyar):

if we plug in these values we get: slope = (60-53)/(1-2) = -7 so slope of function 1 =-7

OpenStudy (priyar):

the second function is in the form of: y=mx+c where m is the slope..here the eq is: y=-8x+60 so, what is m? (just compare this with the general form..u can tell just by looking!) after obtaining both slopes compare them to find the correct option.. Hope this helps!